5. A radio transmitter transmits a signal of 1MHz range. Which are the five possible selections of capacitor and inductor for capture this signal?
1
Expert's answer
2019-11-22T11:16:06-0500

A radio receiver receives at a resonant frequency. The resonant frequency in terms of capacity and inductance is


"f_r=\frac{1}{2\pi\sqrt{LC}}."

Hence:


"L=\frac{1}{4\pi^2f_r^2C}=\frac{2.533\cdot10^{-14}}{C}."

Use this relation:

"C=350\text{ pF},\space L=7.237\cdot10^{-5}\text{ H};\\
C=350\text{ nF},\space L=7.237\cdot10^{-8}\text{ H};\\
C=350\space\mu\text{F},\space L=7.237\cdot10^{-11}\text{ H};\\
C=2\text{ pF},\space L=0.01267\text{ H};\\
C=5\text{ pF},\space L=5.066\cdot10^{-3}\text{ H}."